A=1/2H (B+6) (FOR B)
Answer by stanbon(75887)

A=1/2H (B+6) (FOR B)
Multiply both sides by 2 to get:
2A = H(B+6)
Divide by H to get:
2A/H = B+6
Subtract 6 from both sides to ge:
B= [2A/H]-6
-------------
cheers,
Stan H.